# oMEGACat II - Photometry and proper motions for 1.4 million stars in Omega Centauri and its rotation in the plane of the sky

### v1

@@ 9.4 Public examples on catalog usage
Together with the data products we publish a jupyter notebook [^13] Again this notebook will be made publicly accesible, likely on github, when the paper is accepted for publication by the Journal. For the refereeing process we provide a PDF Export of the notebook that can be used as starting point for usage of our catalog. This example includes:
@@ 10 Conclusions
As a first science result, we determined the plane-of-sky rotation curve of $\omega$ Cen with unprecedented resolution using a combination of our relative proper motions and the absolute proper motions from Gaia. In addition, we obtain a precise measurement of $\omega$ Cen’s inclination of $i=(43.6\pm 1.5)^{\circ}$.
The combined oMEGACat catalogs are already enabling a broad range of interesting science. Ongoing projects are the study of the Age-Metallicity relation of $\omega$ Cen (Clontz et al. in prep.), the automated separation of subpopulations based on photometry and metallicity (Clontz et al. in prep), the discovery of fast-moving stars indicative of an intermediate-mass black hole (Häberle et al. under review), the search for spatial differences in the metallicity distribution (Nitschai et al. in prep.) and the extraction of individual abundances using stacked spectra (Di Stefano et al. in prep.). The MUSE data of the centermost region is well matched to the depth of our proper motion catalog and has revealed a counter-rotating core in the centermost (Pechetti et al., 2024) region. We plan to use the combined dataset to create a dynamical model of this region (Pechetti et al.) and eventually the whole cluster.
A.B. acknowledges support from STScI grant GO-15857. A.B. C.C.,M.A.C., and A.C.S. acknowledge support from STScI grant GO-16777. Based on archival and new observations with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS 5-26555. SK acknowledges funding from UKRI in the form of a Future Leaders Fellowship (grant no. MR/T022868/1). A.F.K. acknowledges funding from the Austrian Science Fund (FWF) [grant DOI 10.55776/ESP542]. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. Gaia, HST astropy, matplotlib, numpy, scipy, hst1pass
@@ Appendix B Photometric Corrections
Appendix C Search for global trends within the proper motion catalog
@@ Appendix D Comparisons with literature catalogs
Figure 24: Similar to Figures 20 - 23, but in this figure we compare the corrected relative proper motions from this work with the absolute proper motions from the Gaia FPR. In comparison with Gaia DR3 the completeness in the center is much better and astrometric residuals are lower. Just as in Figure 23 one can see the clear imprint of the cluster’s rotation, which is further studied in Section 8 and Figure 14.

### v2

@@ 9.4 Public examples on catalog usage
Together with the data products we publish an IPython notebook  [^11] Again this notebook is made publicly accessible in a Zenodo repository (doi:10.5281/zenodo.11104046) that can be used as starting point for the usage of our catalog. The notebook includes:
@@
  • An exemplary calculation on how to propagate the stellar motions from the new catalog to any given epoch while properly accounting for the absolute motion of the cluster and the relative motion of the individual stars
@@ 10 Conclusions
As a first science result, we determined the plane-of-sky rotation curve of $\omega$ Cen with unprecedented resolution using a combination of our relative proper motions and the absolute proper motions from Gaia. In addition, we obtain a precise measurement of $\omega$ Cen’s inclination of $i=(43.9\pm 1.3)^{\circ}$.
The combined oMEGACat catalogs are already enabling a broad range of interesting science. Ongoing projects are the study of the Age-Metallicity relation of $\omega$ Cen (Clontz et al. in prep.), the automated separation of subpopulations based on photometry and metallicity (Clontz et al. in prep), the discovery of fast-moving stars indicative of an intermediate-mass black hole (Häberle et al., 2024), the search for spatial differences in the metallicity distribution (Nitschai et al., 2024) and the extraction of individual abundances using stacked spectra (Di Stefano et al. in prep.). The MUSE data of the centermost region is well matched to the depth of our proper motion catalog and has revealed a counter-rotating core in the centermost (Pechetti et al., 2024) region. We plan to use the combined dataset to create a dynamical model of this region (Pechetti et al.) and eventually the whole cluster.
A.B. acknowledges support from STScI grant GO-15857. A.B. C.C.,M.A.C., and A.C.S. acknowledge support from STScI grant GO-16777. Based on archival and new observations with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS 5-26555. SK acknowledges funding from UKRI in the form of a Future Leaders Fellowship (grant no. MR/T022868/1). A.F.K. acknowledges funding from the Austrian Science Fund (FWF) [grant DOI 10.55776/ESP542]. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. Gaia, HST astropy (Astropy Collaboration et al., 2022), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), IPhyon (Perez & Granger, 2007), hst1pass (Anderson, 2022)
@@ Appendix A Dataset
Tables A and A show show information on all individual exposures used for the creation of our catalog. A compilation of the full data set is also archived at doi:http://dx.doi.org/10.17909/26qj-g090 (catalog 10.17909/26qj-g090).
@@ Appendix B Photometric Corrections
Figures 16 and 17 show maps of the spatially variable photometric corrections derived in Section 6.5. These corrections are a superposition of differential reddening, which has a physical origin, and instrumental effects and zero-point variations.
@@
Appendix C Catalog validation: Search for systematic effects in magnitude and color
In addition to the local trends in magnitude, we also searched for global trends in magnitude and color, by dividing the proper motions in $m_{\rm F625W}$-magnitude, or $m_{\rm F625W}-m_{\rm F814W}$ color bins (see Figure 18). We calculated the median of the two proper-motion components in each bin and did not see any significant deviation from zero, neither for the raw nor for the a-posteriori corrected proper motions (see Figure 18).
@@ Appendix D Comparisons with literature catalogs
D.1 Crossmatch and direct comparison with Bellini et al. (2017a)
We transformed the pixel-based coordinate system of Bellini et al. (2017a) to our own reference system using 6 parameter linear transformations. After that, we used a matching radius of 1 WFC/UVIS pixel (40 mas) to crossmatch stars. This radius is large enough, as the reference epochs of the two catalogs are similar (Bellini et al.: 2007.0; This work: 2012.0). In 5 years the stars will have an RMS displacement of just 0.08 pixel. As it can be seen in Figure 19, the Bellini et al. catalog is fully contained within our larger field and most stars can be crossmatched (465 362/478 477 for the photometric and 242934/245443 for the astrometric catalog).
While the Bellini et al. catalog contains photometry for 26 filters, the focus of our study was on the 6 (3 ACS/WFC, 3 WFC3/UVIS) filters for which we have full coverage out to the half-light radius and the WFC3/UVIS F606W filter for which we have the largest number of photometric measurements in the center. Therefore, there are only 4 filters that we can compare. We show the difference of the method 1 (see Section 4.6) photometry for these 4 filters in Figure 19. The overall agreement is good and only very small systematic shifts of the zero point can be observed (F275W: -0.007; F336W: 0.014; F606W: 0.027; F814W: 0.006). We attribute those differences to the updated instrumental zero point values (see Section 6.1) and the slightly different radii used to create the aperture-photometry-based reference systems. They are of the same order of magnitude as the reported uncertainties on the absolute flux calibration (Calamida et al., 2022).
To directly compare the proper motions, we restrict our analysis to stars brighter than $m_{\rm F606W}=18$, as their statistical errors are lower and potential systematic effects are easier to detect. Figure 20 shows the comparison of the raw proper motions. While the overall agreement is good (RMS of difference $\sim$0.09 mas yr-1 in both components), one can see some low spatial frequency effects with amplitudes of up to $\sim$0.1 mas yr-1. These systematic deviations cannot be explained using the individual proper-motion errors alone (the error distributions are 2.31/2.56 times wider than what would be expected from the proper motions errors alone). This is not unexpected for the raw proper motions and can be attributed to CTE effects and residual distortion and is also why we employed the local a-posteriori corrections (see Section 5.3).
In Figure 21, we compare the proper motions after the local a-posteriori corrections have been applied in both catalogs. As expected, the low spatial frequency pattern has disappeared. Instead, we now can see some granularity which is most likely caused by the spatial scale of the local corrections and the limited number of available reference stars. The errors of the corrected proper motions do account for this additional uncertainty and, therefore, the distribution of the residuals is now much more compatible with the errors in the proper-motion catalogs (1.36 / 1.33 times wider than what would be expected from the proper motions errors). The magnitude dependence of the deviations between the proper-motions from the two catalogs is compared in Figure 22, as expected the deviations increase for fainter stars.
@@
D.2 Crossmatch and direct comparison with Gaia DR3 & FPR
As already described above, the Gaia DR3 is affected by crowding and has been extended by the Gaia FPR. Here, we crossmatch the combined DR3 + FPR Gaia catalog with our new HST catalog. Using a simple geometric cutoff of 160 mas, we find 373 291 stars in common with our catalog (27 123 from DR3 and 346 168 from the FPR). When comparing the completeness of the two catalogs with respect to each other, see Figure 23, one can see very good agreement. When using our new HST catalog as reference, the combined Gaia DR3 + FPR catalog shows good completeness for magnitudes brighter than $m_{\rm F625W}\sim 19$, for fainter stars the Gaia completeness drops sharply. On the other hand, if one uses the Gaia catalog as a reference and restricts it to the region covered by our HST catalog, almost all ($>99\%$) Gaia sources can be recovered in the HST catalog over the full magnitude range. However, the sources brighter than Gaia G14.514.5~{}14.514.5 typically have no HST proper motion measurements, due to saturation in the HST data.
As many of the crossmatched Gaia sources have limited accuracy and are affected by crowding, for the following analysis we limited ourselves to a subset of well-measured stars. We used a simple proper motion error cut of 0.6 mas yr-1 in both components and both datasets. In addition, we restrict the matching radius to 40 mas (1 WFC3/UVIS pixel) for the high-quality subset and only use stars with a minimum proper-motion baseline of 10 years. This high-quality (HQ) subset of crossmatched stars contains 5 897 entries from DR3 and 24 467 from the FPR. In Figure 13 (left), it can be seen that, while the actual stellar density increases towards the center, the number of well-measured Gaia stars decreases. Also, it can be seen in the color-magnitude diagram in Figure 13 (middle panel) that the overlap between the catalogs is limited to a relatively small magnitude range of mostly evolved stars. In the right panel of Figure 13, we study the positional residuals between our catalog and well-measured stars from the Gaia catalogs. Given the different reference epochs of the catalogs (our catalog: 2012.0, Gaia DR3: 2016.0; Gaia FPR: 2017.5), and the fact that we anchored our (co-moving) astrometric reference system on observations from 2002.5, we expect both a systematic shift (based on the absolute cluster motion with respect to 2002.5), and random deviations (caused by the individual stars random motion). As it can be seen in the figure, the absolute deviation is around 101 mas for DR3, and 112 mas for FPR, which is compatible with what we expect from the known absolute cluster motion from our initial 2002.5 epoch towards the respective Gaia reference epochs. The random position deviation has an RMS of $\sim$ 3 mas (DR3) and 4 mas (FPR), also compatible with the displacement expected from the velocity dispersion of the cluster.
Figure 24 (DR3) and 25 (FPR) show the direct comparison and the differences between the proper motion components between the Gaia datasets and our new catalog. While there is an overall good agreement in terms of pure proper-motion values after the absolute proper motion of $\omega$ Cen is accounted for, there is one large-scale systematic trend that leads to differences of up to $\sim$0.3 mas yr-1. The reason for these systematic differences lies in the different approaches used to measure proper motions: While Gaia measures absolute proper motions anchored to a fixed reference frame, we measure our HST proper motions relative to the bulk motion of cluster stars. We further discuss these systematic differences between the two catalogs in the following Section 8, where we use them to obtain a new, accurate measurement of $\omega$ Cen’s rotation curve.
Figure 23: The upper panels show histograms of the magnitude distribution of both our new oMEGACat HST catalog and the combined Gaia DR3+FPR catalog, and the respective cross-matches between them. The ratio of recovered stars over all stars gives us the relative completeness between the catalogs, which is shown in the lower panels. The combined Gaia catalog shows high completeness ($>90\%$) with respect to HST until it sharply drops at faint magnitudes. A 50% level of completeness is reached around $m_{\rm F625W}\sim 21$. The HST catalog is complete ($>99\%$) with respect to Gaia over the full Gaia magnitude range, however at bright magnitudes (Gaia Gmag $<14.5$) typically no HST proper motions are available due to saturation in the HST images.
@@ D.2 Crossmatch and direct comparison with Gaia DR3 & FPR
Figure 25: Similar to Figure 24, but in this figure we compare the corrected relative proper motions from this work with the absolute proper motions from the Gaia FPR. In comparison with Gaia DR3 the completeness in the center is much better and astrometric residuals are lower. Just as in Figure 24 one can see the clear imprint of the cluster’s rotation, which is further studied in Section 8 and Figure 14.
@@ Appendix E Description of Columns in the Data Product Tables
Table E shows the content of the astrometric catalog with explanations for each individual column (see also Section 9.1). Table E shows the content of each of our seven photometric catalogs with explanations for each column (see also Section 9.2). The tables can be downloaded from Zenodo: doi:10.5281/zenodo.11104046.
@@
Appendix F Numerical values of rotation profile
Table 10 shows the numerical values for the plane-of-sky rotation profile determined in Section 8.1. It is also available in machine-readable form in the Zenodo Repository: doi:10.5281/zenodo.11104046
Table 10: Numerical values of the rotation profile from Fig. 14
Lower limit Median radius Upper limit Number of Stars Median tangential Inferred
of bin of stars in bin of bin proper motion Rotation Velocity
[arcsec] [arcsec] [arcsec] - [mas yr-1] [km s-1]
0.0 20.66 30.0 270 0.0548$\pm$0.0324 1.41$\pm$0.83
30.0 47.6 60.0 1020 0.086$\pm$0.0181 2.21$\pm$0.47
60.0 76.65 90.0 1620 0.1478$\pm$0.0131 3.8$\pm$0.34
90.0 105.91 120.0 2423 0.2088$\pm$0.0105 5.38$\pm$0.27
120.0 135.99 150.0 3031 0.2556$\pm$0.0093 6.58$\pm$0.24
150.0 165.6 180.0 3882 0.2562$\pm$0.0083 6.59$\pm$0.21
180.0 195.11 210.0 4361 0.2678$\pm$0.008 6.89$\pm$0.21
210.0 225.38 240.0 4470 0.2648$\pm$0.0078 6.82$\pm$0.20
240.0 254.06 270.0 4186 0.2797$\pm$0.0083 7.2$\pm$0.21
270.0 282.86 300.0 3020 0.2619$\pm$0.0103 6.74$\pm$0.27
300.0 309.92 330.0 985 0.2658$\pm$0.0176 6.84$\pm$0.45
330.0 340.91 360.0 246 0.2455$\pm$0.0421 6.32$\pm$1.08